# Breaking Information Cocoons: A Hyperbolic Graph-LLM Framework for Exploration and Exploitation in Recommender Systems

### v2

@@ Appendix A Proof
In this section, we provide a rigorous mathematical proof of Proposition 4.1, demonstrating the unique advantage of the aligning mechanism in the hyperbolic space.
@@
Appendix B Profile Decoder
@@
To further evaluate the effectiveness of hyperbolic alignment and the expressiveness of our learned hyperbolic embeddings, we introduce a profile decoder designed to reconstruct the original profile from these embeddings, thereby broadening their application beyond recommendation tasks. Our method begins by mapping the hyperbolic embeddings back into Euclidean space to facilitate alignment with traditional network layers. These Euclidean-mapped embeddings are then processed through a mixture of experts (MoE) module. The MoE serves to refine the embeddings, ensuring consistency in both dimensionality and semantic representation (Ma et al., 2024). Once processed, these embeddings are fed into the language model, specifically after the positional layer, to incorporate them within a sequential architecture. To enhance training efficiency and minimize computational cost, we freeze the language model’s parameters, making the MoE module the only trainable component in the architecture.
@@
B.1 Decoder Quality
To assess the effectiveness of our decoder mechanism, we compare the embeddings from our model against those from Euclidean embeddings (LightGCN) and Hyperbolic embeddings (HICF) as inputs to the decoder. The results are shown in Table 5. We use BERTScore (Zhang et al., 2019) to evaluate the semantic alignment of generated sentences with the original profile. Results indicate that our HERec outperforms both Euclidean and hyperbolic baselines, demonstrating that our model’s embeddings support recommendation and retain user/item profile information. This finding underscores that our model captures collaborative information alongside recoverable language semantics: a capability not achieved by previous models.
B.2 Decoder Case Study
@@
 Original item profile:
Ceol Irish Pub in Reno is a popular bar and pub known for its happy hour deals, great service, and friendly atmosphere. It is suitable for those who enjoy trivia nights, live music, and a wide selection of drinks, particularly whiskey. The pub has a welcoming environment, with bartenders who are attentive and quick to serve.
@@
Appendix C Hierarchical Structure Case Study
In this section, we present a case study examining the exploration-exploitation trade-off in the hyperbolic hierarchical structure. We randomly select a user and identify items that share different layers of the lowest common ancestor (LCA) with the user in the hierarchy tree. Ideally, a larger LCA layer indicates that the user and item were grouped earlier in the hierarchy, suggesting a stronger likelihood of interaction. Conversely, a smaller LCA layer implies that the user and item were grouped only after multiple iterations, indicating a weaker likelihood of interaction.
 User Profile:
Item 1 Profile (LCA=10):
Item 2 Profile (LCA=5):

### v3

@@ Appendix A Proof
In this section, we provide a rigorous mathematical proof of Proposition 1, demonstrating the unique advantage of the aligning mechanism in the hyperbolic space.
@@
A.3 Error Bound Analysis
The approximations used in the proof rely on the assumption that $|x|$ and $|y|$ are large, which is common in high-dimensional hyperbolic embedding spaces. In practice, with embedding dimensions such as 50 (as used throughout our experiments), the norms tend to be sufficiently large for the approximation to hold effectively. To further substantiate this assumption, we calculated the error bound as follows:
The exact Lorentz inner product is given by:
$z=\sqrt{1+|x|^{2}}\sqrt{1+|y|^{2}}-|x||y|\cos\theta,$
and is approximated as:
$z_{\text{approx}}=|x||y|(1-\cos\theta).$
Using a Taylor expansion, we have:
$\sqrt{1+|x|^{2}}=|x|\left(1+\frac{1}{2|x|^{2}}-\frac{1}{8|x|^{4}}+\cdots\right).$
Thus, the absolute error becomes:
$|z-z_{\text{approx}}|=\frac{|y|}{2|x|}+\frac{|x|}{2|y|}.$
The relative error in the Lorentz product is bounded by:
$\epsilon_{z}\leq\frac{1}{2}\left(\frac{1}{|x|^{2}}+\frac{1}{|y|^{2}}\right)\frac{1}{1-\cos\theta}.$
Additionally, the approximation error in $\sqrt{z^{2}-1}\approx z$ is given by:
$\sqrt{z^{2}-1}=z\left(1-\frac{1}{2z^{2}}+\cdots\right),$
which implies:
Relative $\text{Relative error}\leq\frac{1}{2z^{2}}.$
Combining these results, the overall error bound is:
Relative Error in |$\text{Relative Error in }|\nabla d_{H}(x,y)|\leq\frac{1}{2|x|^{2}(1-\cos\theta)}+\frac{1}{2z^{2}}.$
For instance, with $|x|=|y|=5$ and $\theta=\frac{\pi}{2}$, the total error is approximately $2.1\%$, demonstrating high approximation accuracy in our setting.
Appendix B Experimental Settings
In this section, we present details of our experimental settings, the distribution of head and tail items (Table 3), implementation details, and an explanation of the diversity metrics used. Notably, the high proportion of tail users (T80) in these datasets highlights the prevalence of long-tail distributions, reaffirming the assumption of a power-law distribution commonly observed in real-world data.
B.1 Implementation Details
We use an embedding dimension of 50 and a training batch size of 1024. The embeddings in hyperbolic space are optimized using Riemannian SGD (RSGD) [2, 25], while the MLP adapter is trained with the Adam optimizer [13]. Training employs early stopping based on Recall@10, with a patience of 10 epochs. User and item profiles are generated using GPT-3.5-turbo and encoded with text-embedding-ada-002 [22] for the Amazon-Books and Yelp datasets, and BERT [8] for Google-Reviews. All experiments are conducted on NVIDIA A40/A100 GPUs with 40GB of memory.
For the final experiments, we perform a grid search over commonly used hyperparameters: learning rate $\in\{1\mathrm{e}{-3},5\mathrm{e}{-3},1\mathrm{e}{-2},5\mathrm{e}{-2},1\mathrm{e}{-1}\}$, weight decay $\in\{1\mathrm{e}{-3},1\mathrm{e}{-2},1\mathrm{e}{-1}\}$, number of negative samples $\in\{10,20,50\}$, and alignment weight $\in\{1\mathrm{e}{-3},1\mathrm{e}{-2},1\mathrm{e}{-1}\}$. A full list of tuned hyperparameters is available in our attached GitHub repository.
Table 3: Statistics of the experimental data.
Dataset #User #Item #Interactions Density
All H20(%) T80(%)
[…]
B.2 Diversity Metrics
$H=-\sum_{i\in\text{set}(\sum_{u}R_{u})}p(i)\log p(i)$ (24)
Appendix C Supplement to Hyperbolic Hierarchical Clustering
In this section, we provide details of the hyperbolic hierarchical clustering algorithm, together with complexity analysis and a detailed case study.
C.1 Hyperbolic Hierarchical Clustering
Algorithm 1 describes the hyperbolic hierarchical clustering procedure used to organize hyperbolic embeddings into a multi-layer structure. The algorithm takes as input the hyperbolic embeddings $X$ and a layer proportion parameter $k$, which controls the reduction rate of points per hierarchical layer.
The maximum number of layers $L$ is set to $\log_{k}(|X|)$, ensuring that each subsequent layer contains approximately $1/k$ times the points of the previous layer. Starting from the bottom layer containing all points, the algorithm iteratively constructs higher layers in a bottom-up fashion.
At each layer $l$, cluster centroids are randomly initialized in hyperbolic space, with the number of clusters equal to $|D_{l}|/k$, where $|D_{l}|$ is the number of points in the current layer. Clustering proceeds by repeatedly assigning each point to its nearest centroid (in terms of hyperbolic distance) and updating each centroid to the hyperbolic mean of its assigned points. This process continues until the centroids converge.
Upon convergence, the centroids from the current layer serve as the points for the next upper layer. By repeating this process, the algorithm constructs a hierarchy of clusters, effectively capturing both the geometric structure and hierarchical relationships inherent in hyperbolic embeddings.
Algorithm 1 Hyperbolic Hierarchical Clustering
  Input: Hyperbolic embeddings $X$, layer proportion $k$
  Set the maximum number of layers $L=\log_{k}(|X|)$
  Initialize the points of bottom layer $D_{L}=X$
  for $l=L$ to $1$ (bottom to top) do
     repeat
        for each point $d_{l,i}\in D_{l}$ do
           Assign $d_{i}$ to the cluster with the nearest centroid:
           $c_{l,i}=\arg\min_{\mu\in M_{l}}d_{\mathcal{H}}(d_{i},\mu)$
        end for
        for each cluster $c_{l,i}\in C_{l}$ do
           $\mu_{l,i}=m_{\mathcal{H}}(\{d\ |\ d\in C_{l,i}\})$
        end for
     until Centroids $M_{l}$ do not change (convergence)
     Initialize the points of the upper layer $D_{l-1}=M_{l}$
  end for
C.2 Complexity Analysis
In this subsection, we analyze the complexity of constructing and applying the hierarchical representation structure. Let $N$ be the total number of users and items, and $k$ be the number of recommendations per user.
Hierarchical Clustering. Building the hierarchy tree involves iterative k-means clustering in hyperbolic space. To optimize efficiency, we assign cluster nodes only to real node embeddings. First, we compute and store pairwise distances, requiring $O(N^{2}/2)$. Each k-means iteration takes $O(i)$, and with $O(\log N)$ levels, the overall complexity of constructing the hierarchy tree is $O(N^{2}/2+i\log N)$.
Recommendation Retrieval & Replacement. Finding the top-$k$ recommendations requires $O(k\log N)$ using nearest neighbour search in hyperbolic space. Adjusting recommendations by retrieving ancestor clusters at layer $l$ requires $O(k)$ time, as it involves only tree traversal.
The proposed hierarchical framework achieves efficient complexity scaling while maintaining flexibility in recommendation adjustments. The hierarchical clustering process, optimized for hyperbolic space, ensures a computationally feasible structure with $O(N^{2}/2+i\log N)$ complexity. The recommendation retrieval and replacement steps remain lightweight, running in $O(k\log N)$ and $O(k)$, respectively. This design enables HERec to dynamically balance exploration and exploitation while ensuring scalability for large-scale recommender systems.
C.3 Hierarchical Structure Case Study
In this subsection, we present a case study examining the exploration-exploitation trade-off in the hyperbolic hierarchical structure. We randomly select a user and identify items that share different layers of the lowest common ancestor (LCA) with the user in the hierarchy tree. Ideally, a larger LCA layer indicates that the user and item were grouped earlier in the hierarchy, suggesting a stronger likelihood of interaction. Conversely, a smaller LCA layer implies that the user and item were grouped only after multiple iterations, indicating a weaker likelihood of interaction.
 User Profile:
Item 1 Profile (LCA=10):
Item 2 Profile (LCA=5):
Data Amazon-books Yelp Google-reviews
Table 5: Standard deviation of key metrics across datasets over 10 independent runs.
Dataset Recall@20 NDCG@20 DIV@20 H@20 EPC@20
Amazon 0.0010 0.0004 0.0067 0.0035 0.0002
Yelp 0.0009 0.0004 0.0091 0.0017 0.0014
Google 0.0010 0.0003 0.0077 0.0008 0.0009
Table 6: Direct message passing within hyperbolic space.
Dataset Recall@20 NDCG@20 DIV@20 H@20 EPC@20
Amazon 0.1721 0.1082 0.5760 12.211 0.8773
Yelp 0.1205 0.0793 0.4005 12.321 0.8755
Google 0.1611 0.0988 0.3532 12.766 0.8731
Appendix D More Experimental Results
In this section, we will show more experimental results to further analyze the flexibility and effectiveness of our HERec.
D.1 Baselines with Feature Enhancement
For further validation, we assessed the impact of feature integration on Euclidean baselines. By incorporating semantic embeddings into the Euclidean space—mirroring our approach in Section 4.2—we created enhanced variants (denoted by “+”). As reported in Table 4, these models generally achieve higher utility scores but do not consistently improve diversity. This reinforces our main findings: while feature integration benefits utility in Euclidean frameworks, it falls short for diversity, highlighting the unique strengths of hyperbolic alignment detailed in Section 4.4.
D.2 Statistical Significance Testing
To ensure the robustness and reliability of our results, we conducted 10 independent runs for each experimental setting and reported the standard deviation for all key metrics. Results are shown in Table 5. Across all datasets and evaluation metrics, we observed consistently low standard deviations, indicating high stability and reproducibility of our findings. This statistical rigor further substantiates the significance of our experimental conclusions.
D.3 Alternative Hyperbolic Message Passing
We also explored an alternative message passing strategy by replacing the original hyperbolic graph collaborative filtering mechanism with direct message passing in hyperbolic space, thereby eliminating the need for tangent space projections. This variant achieves performance comparable to our primary method, results can be found in Table 6. These findings suggest that direct message passing in hyperbolic space is not only a viable alternative but could also serve as a complementary component within our framework, potentially offering additional flexibility for future extensions.
Appendix E Profile Decoder
System Instruction: Describe what types of items this user is likely to enjoy./ Describe what types of users would enjoy this item. Input Prompt: <USER_EMBED>/<ITEM_EMBED> Output: The user would prefer items which …/ The item would be preferred by users who …
@@ Appendix E Profile Decoder
To further evaluate the effectiveness of hyperbolic alignment and the expressiveness of our learned hyperbolic embeddings, we introduce a profile decoder designed to reconstruct the original profile from these embeddings, thereby broadening their application beyond recommendation tasks. Our method begins by mapping the hyperbolic embeddings back into Euclidean space to facilitate alignment with traditional network layers. These Euclidean-mapped embeddings are then processed through a mixture of experts (MoE) module. The MoE serves to refine the embeddings, ensuring consistency in both dimensionality and semantic representation [20]. Once processed, these embeddings are fed into the language model, specifically after the positional layer, to incorporate them within a sequential architecture. To enhance training efficiency and minimize computational cost, we freeze the language model’s parameters, making the MoE module the only trainable component in the architecture.
@@
E.1 Decoder Quality
To assess the effectiveness of our decoder mechanism, we compare the embeddings from our model against those from Euclidean embeddings (LightGCN) and Hyperbolic embeddings (HICF) as inputs to the decoder. The results are shown in Table 7. We use BERTScore [44] to evaluate the semantic alignment of generated sentences with the original profile. Results indicate that our HERec outperforms both Euclidean and hyperbolic baselines, demonstrating that our model’s embeddings support recommendation and retain user/item profile information. This finding underscores that our model captures collaborative information alongside recoverable language semantics: a capability not achieved by previous models.
E.2 Decoder Case Study
@@
Original item profile: Ceol Irish Pub in Reno is a popular bar and pub known for its happy hour deals, great service, and friendly atmosphere. It is suitable for those who enjoy trivia nights, live music, and a wide selection of drinks, particularly whiskey. The pub has a welcoming environment, with bartenders who are attentive and quick to serve.